OpenStudy (anonymous):

Solve. |–4b – 8| + |–1 – b²| + 2b³ ; b = –2

OpenStudy (jdoe0001):

\(\bf |-4\color{red}{b} - 8| + |-1 - \color{red}{b}^2| + 2\color{red}{b}^3 \qquad \qquad b = -2\\\quad \\\quad \\ |-4\color{red}{(2)} - 8| + |-1 - \color{red}{(2)}^2| + 2\color{red}{(2)}^3\)

OpenStudy (jdoe0001):

keep in mind that absolute values are always positive

OpenStudy (anonymous):

/-16/+/-5/+16?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Why is it 2 instead of -2 for the b

OpenStudy (jdoe0001):

ohh shoot... yes... darn... anyhow, you're correct, I missed the " - " sign, missed that, one sec

OpenStudy (jdoe0001):

\(\bf |-4\color{red}{b} - 8| + |-1 - \color{red}{b}^2| + 2\color{red}{b}^3 \qquad \qquad b = -2\\\quad \\\quad \\ |-4\color{red}{(-2)} - 8| + |-1 - \color{red}{(-2)}^2| + 2\color{red}{(-2)}^3\)

OpenStudy (anonymous):

/0/+/3/+16?

OpenStudy (anonymous):

0+3+16=19?

OpenStudy (jdoe0001):

\(\bf |-4b - 8| + |-1 - b^2| + 2b^3 \qquad \qquad b = -2\\\quad \\\quad \\ |-4(-2) - 8| + |-1 - (-2)^2| + 2(-2)^3\\ |0|\\ \qquad \qquad \color{blue}{(-2)^2 \implies -2 \times -2 \implies +2}\\ |-1 - (-2)^2| \implies |-5|\\ \qquad \qquad \color{blue}{(-2)^3 \implies -2 \times -2 \times -2 \implies -2}\\ 2(-2)^3 \implies -16\)